a survey report states that 37% of adults visit their doctor regularly. suppose that 10 adults are selected at random. what is the probability that exactly 8 adults visit their doctor regularly in such group?

Answer by rothauserc(4718)

use the binomial probability formula
:
Probability (P) ( k successes in n trials ) = nCk * p^k * (1-p)^(n-k), where nCk = n! / (k! * (n-k)!)
:
for this problem n=10, k=8, p=0.37
:
P (8 adults visit their doctor regularly out of 10 selected) = 10C8 * (0.37)^8 * (1-0.37)^(10-8) = 0.0063
: